OpenStudy (anonymous):

Choose what the expressions below best represent within the context of the word problem. The length of a rectangle is 12 inches more than its width. What is the width of the rectangle if the perimeter is 42 inches? x best represents . x + 12 best represents .

OpenStudy (texaschic101):

x best represents WIDTH x + 12 best represents LENGTH

OpenStudy (anonymous):

Select the equation that most accurately depicts the word problem. One-half of a certain number is 95. - n = 95 + n = 95 · n = 95 ÷ n = 95